OpenStudy (anonymous):

Read the statements shown below. Statement 1: If it is a triangle, then it has three sides. Statement 2: If it does not have three sides, then it is not a triangle. Are the two statements logically equivalent? No, only one statement is true. Yes, both statements are false. Yes, both statements are true. No, both statements are false.

OpenStudy (anonymous):

Remember what I said earlier? If p, then q. If we negate both sides, we obtain if not p, then not q. This statement is still logically equivalent as you can see it as both sides of an equation receiving the "same treatment".

OpenStudy (anonymous):

would it be: Yes, both statements are true.

OpenStudy (anonymous):

Well, do airplanes fly?

OpenStudy (maheshmeghwal9):

(p+q)=-(p+q)=-p-q yes both r true:)

OpenStudy (anonymous):

hahaha yes @alanli123

OpenStudy (anonymous):

thank you @maheshmeghwal9 & @alanli123 <3

OpenStudy (maheshmeghwal9):

yw:)